# Berkeley, CA 94720, USA

### v2

@@ 4.1 Ray-Singer Torsion for $\Gamma=\mathbb{Z}_{k}$
$H^{i}(S^{2n+1}/\mathbb{Z}_{k},\mathbb{Z})=\left\{\begin{array}[]{ll}\mathbb{Z}&\text{for $i=0,2n+1$,}\\ \mathbb{Z}_{k}&\text{for $i=2,4,\dots,2n$,}\\ 0&\text{for $i=1,3,\dots,2n-1$}.\\ \end{array}\right.$ (45)
@@
where $\rho$ counts inequivalent non-trivial representations of $\Gamma$ [^4] The factor $\dim(\rho)$ does not appear in Friedmann:2002ty , because only one-dimensional representations are needed there, for $\Gamma=\mathbb{Z}_{k}$. The factor $\log\tau(S^{3}/\Gamma,\rho)$ has been originally computed by Ray in ray1970reidemeister . Ray gave a formula for the torsion for any orbifold $S^{3}/\Gamma$ as long as $\Gamma$ acts freely. An improved approach was used in Dowker:2009in , where an effective tau function is used to simplify Ray’s derivation. To describe it, first define the representation $\rho_{m}:\,\mathbb{Z}_{k}\rightarrow\text{U}(1)$ where the generator $\mu$ of $\mathbb{Z}_{k}$ is represented by
@@
with representations $\rho_{m}$ labeled by an integer $m$. In the range $m=1,\dots,k-1$, we get non-trivial inequivalent representations of $\mathbb{Z}_{k}$. Using the approach of Dowker:2009in , the formula for the analytic torsion on a Lens space with a non-trivial representation is then
@@
Here, $p$ labels conjugacy classes of $\Gamma$, and $l$ runs over non-trivial ones. For a complete derivation we refer to Dowker:2009in . Summing over $p$ fixes $l\equiv m\mod k$, leaving the formula for torsion for $\rho_{m}$:
@@
As a quick check on formula (47), we can sum over all non-trivial representations to get
@@
All our fields are in the adjoint representation $\lambda$ of the gauge group SU$(2)$, and we need to decompose $\lambda$ into irreducible (one-dimensional) representations of the subgroup $\Gamma=\mathbb{Z}_{k}$:
@@ 4.2 Ray-Singer Torsion for $\Gamma=\text{Dic}_{k}$
where we also used the fact that the order of Dick is $4k$. Interestingly, this result can also be obtained directly from the results for Lens spaces, using the fact that Dick, as a set, can be represented as the union of certain cyclic subgroups. The log of the Ray-Singer torsion computes the regularized sum of the logs of eigenvalues of harmonic functions on $S^{3}$ that are invariant under $\Gamma$, and this sum can be recast in terms of sums of logs of eigenvalues of harmonic functions that are invariant only under subgroups of $\Gamma$, with added signs to achieve cancellations so that overall only eigenvalues of completely invariant harmonic functions will remain. Specifically, we can use the observation that Dick can be represented as the union of a cyclic subgroup isomorphic to $\mathbb{Z}_{2k}$ and $k$ cyclic subgroups isomorphic to $\mathbb{Z}_{4}$ ikeda1995spectrum . One needs to take into account the fact that some of their intersections are non-trivial and are isomorphic to $\mathbb{Z}_{2}$ subgroups. Overall, one finds ikeda1995spectrum ; Dowker:2009in :
@@
For non-trivial representations of Dick, the calculation will be different depending on whether the representations are one- or two-dimensional. Again, we follow the method described in Dowker:2009in . We start the discussion with conjugacy classes of Dick; there are $k+3$ of them, and using the generators $(r,s)$ they can be written as:
@@
where $m$ runs from 00 to $12k-$ 1. Excluding the trivial one, the first $k$ conjugacy classes correspond to a rotation by an angle $\pi l/k$ where $l=1,\dots,k$. The last two are rotations by $\pi/2$ for all values of $m$. The torsion for for a representation $\rho$ was explicitly computed in Dowker:2009in , and we will now quote their result. Using $p$ to label the conjugacy classes (55), starting from $p=0$, and using $\theta(p)$ to denote the rotation angles that we just described above, the formula of Dowker:2009in is
$\log\tau(S^{3}/\text{Dic}_{k},R)=\frac{1}{4k}\sum_{p=0}^{k+2}\chi_{\rho}(p)c_{p}\sum_{q=1}^{2k-1}2\cos[q\theta(p)]\log\left[2\sin{\left(\frac{\pi q}{2k}\right)}\right],$ (56)
where $c_{p}$ is the size of the $p^{th}$ conjugacy class.
@@
for $m=1,3,\dots$ not exceeding $k-1$. Therefore, we can restrict the first sum of (56), the one over $p$, to only the first $k$ conjugacy classes of (55). In this way, (56) simplifies to
@@
Below, we list the numerical values for the first few $k$’s. We also confirm results given in Dowker:2009in , where cases $k=2,3,4$ were presented. For completion, we start first with results for one dimensional representations and then we follow with the two dimensional ones.
@@
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{4}}}})=2,$ (64)
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\overline{\mathbf{1}}_{{{3}}}})=4.$ (65)
@@
therefore the adjoint action gives the same decomposition as the one we encountered before, $\gamma\rightarrow\omega^{2}\oplus\mathbf{1}\oplus\omega^{-2}$. Then, using the representations from (5), we obtain for even $k$
$\tau(S^{3}/\text{Dic}_{k},\{{\mathbf{1}_{{{2}}}},{\mathbf{1}_{{{3}}}},{\mathbf{1}_{{{4}}}}\})=\tau(S^{3}/\text{Dic}_{k},\mathbf{1})^{3},$ (66)
@@
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{2}}}})$ $=\tau(S^{3}/\text{Dic}_{k},\mathbf{1})^{3},$ (67)
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{2}}}})^{2}\tau(S^{3}/\text{Dic}_{k},\mathbf{1}).$ (68)
@@
$\gamma\rightarrow{\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}},$ (69)
@@
$\tau(S^{3}/\text{Dic}_{k},\gamma)=\tau(S^{3}/\text{Dic}_{k},{\mathbf{2}_{{2m}}})\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{2}}}}).$ (70)
@@ 5 Hilbert Spaces of the (2,0)-theory on Spherical Orbifolds
We can now combine the results of the Ray-Singer torsion for various flat connections to explicitly compute the Witten indices of spaces of the $(2,0)$-theory on $X_{5}=S^{3}/\Gamma\times S^{2}$, by evaluating (38). First we consider only gauge bundles with trivial Stiefel-Whitney classes. Later on, we extend the discussion to non-trivial Stiefel-Whitney classes as well. As we have seen in §3, because of the presence of $\rho$ and $\chi$ zero modes for all the abelian flat connections, only non-abelian flat connections contribute to the sum over $A$ in (38). For $\Gamma=\mathbb{Z}_{k}$, the Witten index therefore vanishes, because all flat connections are abelian. Thus, we restrict our attention to $\Gamma=\text{Dic}_{k}$. The non-abelian flat connections are in one-to-one correspondence with irreducible $2$-dimensional representations $R_{m}={\mathbf{2}_{{1}}},{\mathbf{2}_{{3}}},\dots$ of $\text{Dic}_{k}$, and since our fields transform in the adjoint representation we find, using (7),
$Z(S^{3}/\text{Dic}_{k}\times S^{2})=\sum_{\begin{array}[]{c}\text{irrep $R_{m}$}\\ \dim R_{m}=2\\ \end{array}}\frac{1}{2}\tau(\text{Dic}_{k},R_{2m}\oplus{\mathbf{1}_{{{2}}}}).$ (74)
where we write $\tau(\Gamma,R_{2m}\oplus{\mathbf{1}_{{{2}}}})$ as a shorthand for $\tau(S^{3}/\Gamma,A[R_{2m}\oplus{\mathbf{1}_{{{2}}}}])$, with $A[R_{2m}\oplus{\mathbf{1}_{{{2}}}}]$ being a flat connection whose holonomy corresponds to the adjoint representation $R_{2m}\oplus{\mathbf{1}_{{{2}}}}$. Combining the results of §4.2, we find
$Z(S^{3}/\text{Dic}_{k}\times S^{2})=\left\{\begin{array}[]{ll}4&\text{for $k=2$;}\\ 2&\text{for $k\geq 3$.}\\ \end{array}\right.$
The results of §4.2 establish this up to $k=6$, but in (82) below we show it for any $k$.
For a non-trivial Stiefel-Whitney class, based on the discussion at the end of §2, for odd $k$ the corresponding flat connections are all abelian, and hence admit fermionic zero modes and do not contribute to the partition function. For even $k$, we have one non-trivial Stiefel-Whitney class that supports a non-abelian flat connection associated with the representation (12). It can be easily checked that for $k>2$ it gives the same partition function as the trivial Stiefel-Whitney class, and therefore the Witten index of space with ${\mathfrak{v}}=0$ in (10) is double that of (5), and if ${\mathfrak{v}}\neq 0$ the Witten index vanishes due to cancellations between ${\mathfrak{w}}=0$ and ${\mathfrak{w}}\neq 0$ partition functions. For $k=2$, if ${\mathfrak{w}}\neq 0$ we get ${\mathcal{Z}}_{\mathfrak{w}}=0$, and therefore the Witten index (10) is $4$ for all ${\mathfrak{v}}$.
@@ 5.1 Hilbert Spaces with Surface Operator Insertions
We wish to generalize the results of §5 by inserting external sources on $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$. A possible external source for the $(2,0)$-theory comes in the form of a string along a closed one-dimensional curve $C\subset X_{5}$, and to compute the Witten index with the sources, we modify the partition function (74) into a correlation function of BRST invariant surface operators along $C\times S^{1}$, for each of the curves $C$. The surface operators are labeled by irreducible representations ${\widetilde{R}}$ of the gauge group SU$(2)$, and the curves $C$ are characterized by their homotopy, which corresponds to an element $\gamma\in\Gamma$. Thus, we can insert ${\mathfrak{n}}$ timelike surface operators corresponding to elements
@@
of the 5d SYM’s gauge group SU$(2)$. Upon the compactification on $S^{1}$ the surface operators reduce to Wilson loops of the BRST invariant gauge field ${\mathcal{A}}$ in representations ${\widetilde{R}}_{1},\dots,{\widetilde{R}}_{\mathfrak{n}}$. (See also Rey:1998ik ; Maldacena:1998im for BPS Wilson loops.) For simplicity, we will take all the Wilson loops to be in the fundamental representation of SU$(2)$, and set ${\widetilde{R}}_{1}=\cdots={\widetilde{R}}_{\mathfrak{n}}=\mathbf{2}$.
@@
where $\tau(\Gamma,R)$ is the $1$-loop contribution which we computed using the Ray-Singer torsion in §4 and $\chi_{R}(\gamma_{i})$ is the character of the representation $R$ evaluated on the element $\gamma_{i}\in\Gamma$.
The expression $I(C_{1},\dots,C_{\mathfrak{n}})$ is the Witten index of the Hilbert space of the $\text{SU}(2)$ twisted $(2,0)$ theory on $(S^{3}/\Gamma)\times S^{2}$ with surface operator insertions that stretch in time and along $C_{1},\dots,C_{\mathfrak{n}}$. This shows that $I(C_{1},\dots,C_{\mathfrak{n}})$ must be an integer, and below we will argue that this is indeed the case. In general, the characters $\chi_{R}(\gamma_{i})$ appearing in (80) are not necessarily integers, and in fact are given by irrational numbers in many examples, as can be seen in (57). Without loss of generality, we can assume that all $\gamma_{i}$ in (80) are of the form $r^{p_{i}}$, for some integers $p_{i}$, otherwise the characters and the index $I(C_{1},\dots,C_{\mathfrak{n}})$ vanish. Therefore, used the torsion (62), we need to show that
$\frac{2^{{\mathfrak{n}}+3}}{k}\sum_{m=1,3,\dots}^{m\leq k-1}\sin^{2}\left(\frac{\pi m}{k}\right)\cos\left(\frac{\pi mp_{1}}{k}\right)\cos\left(\frac{\pi mp_{2}}{k}\right)\cdots\cos\left(\frac{\pi mp_{{\mathfrak{n}}}}{k}\right)\in\mathbb{Z}.$ (81)
This follows from the statement for ${\mathfrak{n}}=1$,
$\frac{16}{k}\sum_{m=1,3,\dots}^{m\leq k-1}\sin^{2}\left(\frac{\pi m}{k}\right)\cos\left(\frac{\pi mp}{k}\right)=\left\{\begin{array}[]{ll}8&\text{for $k=2$ and $p=0$,}\\ -8&\text{for $k=2$ and $p=k$,}\\ 4&\text{for $k\geq 3$ and $p=0$,}\\ -4&\text{for $k\geq 3$ and $p=k$,}\\ 0&\text{for $k=2$ and $p=1$,}\\ 2&\text{for $k=3$ and $p=1$,}\\ -2&\text{for $k=3$ and $p=2$,}\\ 0&\text{for $k=4$ and $p=1,2,3$,}\\ -2&\text{for $k\geq 5$ and $p=2$,}\\ 2&\text{for $k\geq 5$ and $p=k-2$,}\\ 0&\text{for $k\geq 5$ and $k\geq p\notin\{2,k-2\}$,}\end{array}\right.$ (82)
@@
$\frac{16}{k}\sum_{m=0,2,\dots}^{m\leq k-1}\sin^{2}\left(\frac{\pi m}{k}\right)\cos\left(\frac{\pi mp}{k}\right)=\left\{\begin{array}[]{ll}0&\text{for $k=2$ and $p=0,1,2$,}\\ 4&\text{for $k\geq 4$ and $p=0,k$,}\\ -4&\text{for $k=4$ and $p=2$,}\\ 0&\text{for $k=4$ and $p=1,3$,}\\ -2&\text{for $k\geq 6$ and $p=2$,}\\ -2&\text{for $k\geq 6$ and $p=k-2$,}\\ 0&\text{for $k\geq 6$ and $k\geq p\notin\{2,k-2\}$.}\end{array}\right.$ (83)

### v3

@@ 4.1 Ray-Singer Torsion for $\Gamma=\mathbb{Z}_{k}$
[…]
@@
where $\rho$ counts inequivalent nontrivial representations of $\Gamma$. [^4] The factor $\dim(\rho)$ does not appear in Friedmann:2002ty , because only one-dimensional representations are needed there, for $\Gamma=\mathbb{Z}_{k}$. The factor $\log\tau(S^{3}/\Gamma,\rho)$ has been originally computed by Ray in ray1970reidemeister . Ray gave a formula for the torsion for any orbifold $S^{3}/\Gamma$ as long as $\Gamma$ acts freely. An improved approach was used in Dowker:2009in , where an effective tau function is used to simplify Ray’s derivation. To describe it, first define the representation $\rho_{m}:\,\mathbb{Z}_{k}\rightarrow\text{U}(1)$ where the generator $\mu$ of $\mathbb{Z}_{k}$ is represented by
@@
with representations $\rho_{m}$ labeled by an integer $m$. In the range $m=1,\dots,k-1$, we get nontrivial inequivalent representations of $\mathbb{Z}_{k}$. Using the approach of Dowker:2009in , the formula for the analytic torsion on a Lens space with a nontrivial representation is then
@@
Here, $p$ labels conjugacy classes of $\Gamma$, and $l$ runs over nontrivial ones. For a complete derivation we refer to Dowker:2009in . Summing over $p$ fixes $l\equiv m\mod k$, leaving the formula for torsion for $\rho_{m}$:
@@
As a quick check on formula (51), we can sum over all nontrivial representations to get
@@
All our fields are in the adjoint representation $\lambda$ of the gauge group SU(2)(2), and we need to decompose $\lambda$ into irreducible (one-dimensional) representations of the subgroup $\Gamma=\mathbb{Z}_{k}$:
@@ 4.2 Ray-Singer Torsion for $\Gamma=\text{Dic}_{k}$
where we also used the fact that the order of Dick is $4k$. Interestingly, this result can also be obtained directly from the results for Lens spaces, using the fact that Dick, as a set, can be represented as the union of certain cyclic subgroups. The log of the Ray-Singer torsion computes the regularized sum of the logs of eigenvalues of harmonic functions on $S^{3}$ that are invariant under $\Gamma$, and this sum can be recast in terms of sums of logs of eigenvalues of harmonic functions that are invariant only under subgroups of $\Gamma$, with added signs to achieve cancellations so that overall only eigenvalues of completely invariant harmonic functions will remain. Specifically, we can use the observation that Dick can be represented as the union of a cyclic subgroup isomorphic to $\mathbb{Z}_{2k}$ and $k$ cyclic subgroups isomorphic to $\mathbb{Z}_{4}$ ikeda1995spectrum . One needs to take into account the fact that some of their intersections are nontrivial and are isomorphic to $\mathbb{Z}_{2}$ subgroups. Overall, one finds ikeda1995spectrum ; Dowker:2009in :
@@
For nontrivial representations of Dick, the calculation will be different depending on whether the representations are one- or two-dimensional. Again, we follow the method described in Dowker:2009in . We start the discussion with conjugacy classes of Dick; there are $k+3$ of them, and using the generators $(r,s)$ they can be written as:
@@
where $m$ runs from 0 to 2k12k-1. Excluding the trivial one, the first $k$ conjugacy classes correspond to a rotation by an angle $\pi l/k$ where $l=1,\dots,k$. The last two are rotations by $\pi/2$ for all values of $m$. The torsion for a representation $\rho$ was explicitly computed in Dowker:2009in , and we will now quote their result. Using $p$ to label the conjugacy classes (59), starting from $p=0$, and using $\theta(p)$ to denote the rotation angles that we just described above, the formula of Dowker:2009in is
$\log\tau(S^{3}/\text{Dic}_{k},R)=\frac{1}{4k}\sum_{p=0}^{k+2}\chi_{R}([p])c_{p}\sum_{q=1}^{2k-1}2\cos[q\theta(p)]\log\left[2\sin{\left(\frac{\pi q}{2k}\right)}\right],$ (60)
where $c_{p}$ is the size of the $p^{th}$ conjugacy class, and $\chi_{R}([p])$ is the value of the character of the representation $R$ of $\text{Dic}_{k}$ on any element of the $p^{th}$ conjugacy class of $\text{Dic}_{k}$.
@@
for $m=1,3,\dots$ not exceeding k1k-1. Therefore, we can restrict the first sum of (60), the one over $p$, to only the first $k$ conjugacy classes of (59). In this way, (60) simplifies to
@@
Below, we list the values for the first few $k$’s. We also confirm results given in Dowker:2009in , where cases $k=2,3,4$ were presented. For completion, we start first with results for one dimensional representations and then we follow with the two dimensional ones.
@@
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{4}}}})=2,\qquad(k=2,4,6,\dots),$ (68)
$\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\overline{\mathbf{1}}_{{{3}}}})=4,\qquad(k=3,5,7,\dots).$ (69)
@@
therefore the adjoint action gives the same decomposition as the one we encountered before, $\gamma\cong\operatorname{Sym}^{2}\lambda\rightarrow\omega^{2}\oplus\mathbf{1}\oplus\omega^{-2}$. Then, using the representations from (5), we obtain for even $k$
$\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{1}_{{{2}}}})$ $=\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{1}_{{{3}}}})$
$=\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{1}_{{{4}}}})=\tau(S^{3}/\text{Dic}_{k},\mathbf{1})^{3},$ (70)
@@
$\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{1}_{{{2}}}})$ $=\tau(S^{3}/\text{Dic}_{k},\mathbf{1})^{3},$ (71)
$\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{1}_{{{3}}}})$ $=\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{2}}}})^{2}\tau(S^{3}/\text{Dic}_{k},\mathbf{1}).$ (72)
@@
$\gamma\cong\operatorname{Sym}^{2}{\mathbf{2}_{{m}}}\rightarrow{\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}},$ (73)
@@
$\tau(S^{3}/\text{Dic}_{k},\gamma)$ $=\tau(S^{3}/\text{Dic}_{k},\operatorname{Sym}^{2}{\mathbf{2}_{{m}}})$
$=\tau(S^{3}/\text{Dic}_{k},{\mathbf{2}_{{2m}}})\tau(S^{3}/\text{Dic}_{k},{\mathbf{1}_{{{2}}}})=\frac{16}{k}\sin^{2}\left(\frac{\pi m}{k}\right).$ (74)
@@ 5 Hilbert Spaces of the (2,0)-theory on Spherical Orbifolds
We can now combine the results of the Ray-Singer torsion for various flat connections to explicitly compute the Witten indices of spaces of the $(2,0)$-theory on $X_{5}=S^{3}/\Gamma\times S^{2}$, by evaluating (42). First we consider only gauge bundles with trivial Stiefel-Whitney classes. Later on, we extend the discussion to nontrivial Stiefel-Whitney classes as well. As we have seen in §3, because of the presence of $\rho$ and $\chi$ zero modes for all the abelian flat connections, only non-abelian flat connections contribute [^5] In Bak:2017jwd , an additional fermionic “gauge symmetry” was imposed whereby fermionic fields are shifted by terms proportional to harmonic forms, which allows for nonzero contributions from flat connections with zero modes (see Appendix D.2 of Bak:2017jwd ), but we do not impose this additional symmetry. to the sum over $A$ in (42). For $\Gamma=\mathbb{Z}_{k}$, the Witten index therefore vanishes, because all flat connections are abelian. Thus, we restrict our attention to $\Gamma=\text{Dic}_{k}$. The non-abelian flat connections are in one-to-one correspondence with irreducible $2$-dimensional representations $R_{m}={\mathbf{2}_{{1}}},{\mathbf{2}_{{3}}},\dots$ of $\text{Dic}_{k}$, and since our fields transform in the adjoint representation we find, using (7),
$Z(S^{3}/\text{Dic}_{k}\times S^{2})={\mathcal{C}}\sum_{\begin{array}[]{c}\text{irrep $R_{m}$}\\ \dim R_{m}=2\\ \end{array}}\frac{1}{2}\tau(\text{Dic}_{k},R_{2m}\oplus{\mathbf{1}_{{{2}}}}).$ (78)
where we write $\tau(\Gamma,R_{2m}\oplus{\mathbf{1}_{{{2}}}})$ as a shorthand for $\tau(S^{3}/\Gamma,A[R_{2m}\oplus{\mathbf{1}_{{{2}}}}])$, with $A[R_{2m}\oplus{\mathbf{1}_{{{2}}}}]$ being a flat connection whose holonomy corresponds to the adjoint representation $R_{2m}\oplus{\mathbf{1}_{{{2}}}}$. As we will argue in §6, this partition function can be also computed using Verlinde formula for Chern-Simons theory on $S^{2}$. From this observation we will find that ${\mathcal{C}}=1$. Combining the results of §4.2, from (78) we find
[…]
For a nontrivial Stiefel-Whitney class, based on the discussion at the end of §2, for odd $k$ the corresponding flat connections are all abelian, and hence admit fermionic zero modes and do not contribute to the partition function. For even $k$, we have one nontrivial Stiefel-Whitney class that supports a non-abelian flat connection associated with the representation (16). It can be easily checked that for $k>2$ it gives the same partition function as the trivial Stiefel-Whitney class, and therefore the Witten index of space with ${\mathfrak{v}}=0$ in (10) is double that of (81), and if ${\mathfrak{v}}\neq 0$ the Witten index vanishes due to cancellations between ${\mathfrak{w}}=0$ and ${\mathfrak{w}}\neq 0$ partition functions. For $k=2$, if ${\mathfrak{w}}\neq 0$ we get ${\mathcal{Z}}_{\mathfrak{w}}=0$, and therefore the Witten index (10) is $4$ for all $\mathfrak{v}$}.
@@ 5.1 Hilbert Spaces with Surface Operator Insertions
We wish to generalize the results of §5 by inserting external sources on $X_{5}=S^{3}/\text{Dic}_{k}\times S^{2}$. A possible external source for the $(2,0)$-theory comes in the form of a string along a closed one-dimensional curve $C\subset X_{5}$, and to compute the Witten index with the sources, we modify the partition function (78) into a correlation function of BRST invariant surface operators along $C\times S^{1}$, for each of the curves $C$. The surface operators are labeled by irreducible representations $\widetilde{R}$} of the gauge group SU(2)(2), and the curves $C$ are characterized by their homotopy, which corresponds to an element $\gamma\in\Gamma$. Thus, we can insert $\mathfrak{n}$} timelike surface operators corresponding to elements
@@
of the 5d SYM’s gauge group SU(2)(2). Upon the compactification on $S^{1}$ the surface operators reduce to Wilson loops of the BRST invariant gauge field $\mathcal{A}$} in representations ${\widetilde{R}}_{1},\dots,{\widetilde{R}}_{\mathfrak{n}}$. (See also Rey:1998ik ; Maldacena:1998im for BPS Wilson loops.) For simplicity, we will take all the Wilson loops to be in the fundamental representation of SU(2)(2), and set ${\widetilde{R}}_{1}=\cdots={\widetilde{R}}_{\mathfrak{n}}=\mathbf{2}$.
@@
where $\tau(\Gamma,R)$ is the $1$-loop contribution which we computed using the Ray-Singer torsion in §4 and $\chi_{R}(\gamma_{i})$ is the character of the representation $R$ evaluated on the element $\gamma_{i}\in\Gamma$, we also set ${\mathcal{C}}=1$ based on the discussion above.
The expression $I(C_{1},\dots,C_{\mathfrak{n}})$ is the Witten index of the Hilbert space of the $\text{SU}(2)$ twisted $(2,0)$ theory on $(S^{3}/\Gamma)\times S^{2}$ with surface operator insertions that stretch in time and along $C_{1},\dots,C_{\mathfrak{n}}$. This shows that $I(C_{1},\dots,C_{\mathfrak{n}})$ must be an integer, and below we will argue that this is indeed the case. In general, the characters $\chi_{R}(\gamma_{i})$ appearing in (87) are not necessarily integers, and in fact are given by irrational numbers in many examples, as can be seen in (61). Without loss of generality, we can assume that all $\gamma_{i}$ in (87) are of the form $r^{p_{i}}$, for some integers $p_{i}$, otherwise the characters and the index $I(C_{1},\dots,C_{\mathfrak{n}})$ vanish. Therefore, using the torsion (4.2), we find the Witten index with $\mathfrak{n}$} surface operator insertions labeled by $\pi_{1}(S^{3}/\text{Dic}_{k})$ elements $r^{p_{1}},\cdots,r^{p_{\mathfrak{n}}}$:
$\text{Index}_{k,{\mathfrak{n}}}(p_{1},\cdots,p_{\mathfrak{n}})=$ (88)
$\frac{2^{{\mathfrak{n}}+3}}{k}\sum_{m=1,3,\dots}^{m\leq k-1}\sin^{2}\left(\frac{\pi m}{k}\right)\cos\left(\frac{\pi mp_{1}}{k}\right)\cos\left(\frac{\pi mp_{2}}{k}\right)\cdots\cos\left(\frac{\pi mp_{{\mathfrak{n}}}}{k}\right).$
Here we expanded (87) and replaced $R$ with ${\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}}$. We also used (4.2) to substitute
That (88) is an integer follows from the same statement for ${\mathfrak{n}}=1$,
[…]
@@
[…]